# On the Origin of Abundance Variations in the Milky Way’s High- Plateau

### v1

@@ 4.2 Scatter Around the SNIa/CCSN Correlation
Broadly, most of the more well-measured elements (e.g., Si, Ca, Al, Fe, Ni, Mn) show similar RSE between the data and prediction, indicating that measurement noise is sufficient to explain the remaining variation after contributions from a varying SNIa/CCSN ratio are removed. In other words, the amount of SNIa contribution and observational errors are the primary sources of variation on the plateau for these elements. Many of the elements that have larger observational uncertainties (e.g., S, Na, K, Cr, V, Co, Cu, and Ce) have higher observed RSE than predicted RSE. The explanation may be that the reported measurement uncertainties are still underestimated, or that the errors have non-Gaussian tails, since these elements have relatively weak lines that are often affected by blending. In some cases, especially Ce and perhaps C+N and Na, the larger observed RE could reflect intrinsic scater in an enrichment process that is neither CCSN nor SNIa.
@@ 4.3 Covariances Between Elements
As we are calculating the covariances of precomputed residuals, the expectation value in Equation 10, , is the mean of the residuals over the entire plateau sample which should be close to, but may not be exactly, 0.
Figure 11 shows the covariance matrices for the observed residuals from the median, [X/Mg]med,obs (as in Figures 7-8, and gray points of 9), in the left panel and the covariance of residuals from the two-process model, [X/Mg]2proc, in the right panel. The size of each dot corresponds to the magnitude of the covariance, and the diagonals represent the variance of the residuals in each element. Blue filled dots represent positive covariances while open red dots represent negative covariances.
@@
The left panel of Figure 11 has many positive covariances, indicating that most elements tend to deviate from the median in the same direction. In particular, the Fe-peak elements, highlighted by the yellow square, all have significant positive residuals; this is consistent with our previous result that variation in SNIa/CCSN contribution is a significant source of correlated abundance scatter on the low-Ia plateau. This effect is also seen, though to a lesser extent, in the elements highlighted by the purple box.
@@
The two-process model captures the variation in the SNIa/CCSN ratio, so that the covariance in the two-process residuals should have the covariance from SNIa/CCSN variation removed. The right panel of Figure 11 shows the covariance matrix of the two-process residuals. As expected, the signs of the two-process residual covariances are much more mixed, compared to the mostly-positive median residual covariances. The magnitude of the covariances between Fe-peak elements have significantly decreased in the two-process residuals, and some have even changed sign, compared to the median residuals. However, V, Mn, and Co, which are odd-Z Fe-peak elements, still show some positive correlation with each other in the two-process residuals. The elements already have small covariances, and their magnitudes are not significantly different between the median and two-process residuals. However, the Ca-O and Ca-Si covariances have different signs. Similarly, the light odd-Z elements like C+N, Na, and Al do not show any obvious correlations with each other in their two-process residuals, though some covariance signs change from the median residuals. The two-process residuals of C+N and Na are less correlated with those of the Fe-peak elements than their median residuals, but are still positively correlated with Ce. Since C+N and Na both have contribution from a delayed process, this result indicates that their delayed process is not fully explained by the same delayed process (SNIa) producing the Fe-peak elements; furthermore, as Ce is an AGB element, this suggests that an AGB process also contributes to C+N and Na.
@@
W22 removed outliers with a two-process residual greater than 10 times the measurement error for that element from their covariance calculations. If we also perform this cut, we find similar results, where the most affected element is Ce. More specifically, the Ce-(C+N) intrinsic [X/Mg]2proc covariance and Ce-Al [X/Mg]med,obs covariance change sign from positive to negative, while the Ce-Ni, Ce-Mn, and Ce-Co [X/Mg]med,obs covariances change from negative to positive. The rms scatter of Ce (diagonal element) also decreases. The majority of the removed outliers appear to be the same population of stars with strongly enhanced Ce and C+N found in W22, who propose that they are rare products of AGB mass transfer, where extra C+N and Ce from the AGB star are deposited onto the observed star’s photosphere. Their presence here indicates that such mass transfer events occur in stars on the low-Ia plateau.
@@ 5 Discussion
First, we tested the effect of the assumed Fe-to-Mg ratio from pure CCSN enrichment, [Fe/Mg]cc, on the two-process model as formalized by W22. Our key result is that when the process amplitudes and are inferred from only [Mg/H] and [Fe/Mg], the predicted linear {X/Mg} abundance is simply a linear interpolation (or extrapolation) of the median {X/Mg} ratios in the high- and low-Ia populations along {Fe/Mg}. This result was not obvious from previous works because they used logarithmic abundance ratios [X/Mg] instead of linear ratios {X/Mg}. It directly follows from this result that the predicted abundances are independent of the assumed [Fe/Mg]cc, and we show that this independence holds to an excellent approximation even when using a multi-element fit for and .
@@
Our results clearly establish that there is a systematic difference in the SNIa/CCSN enrichment ratio between stars at the “top” of the low-Ia plateau and stars at the “bottom,” a range of roughly 0.1 dex in [Mg/Fe]. This rules out the idea that the median [Mg/Fe] on the low-Ia plateau represents the core collapse yield ratio [Mg/Fe]cc, with abundance ratios scattering stochastically above and below because of, e.g., stochastic sampling of the CCSN population. It remains possible that [Mg/Fe]cc lies at the upper boundary of the plateau, at [Mg/Fe] (See Figure 1). However, our results also lend credence to the idea that the true CCSN ratio is substantially higher, perhaps [Mg/Fe], and the observed plateau in the MW thick disk already includes substantial SNIa enrichment (Maoz & Graur, 2017; Conroy et al., 2022). Maintaining a nearly flat plateau at this intermediate [Mg/Fe] requires accelerating star formation if the ISM is well mixed (Conroy et al., 2022; Chen et al., 2023), though a similar effect may be achieved if early enrichment is spread throughout the circumgalactic medium and only slowly returns to the star-forming ISM (Mason et al., 2024).
@@
For most elements we find good agreement between the predicted and observed slope of [X/Mg]med vs. , but there are exceptions. Most notably, the observed slope is steeper than predicted for C+N and shallower than predicted for Ce. For both abundances, we expect the delayed contribution comes from AGB stars rather than SNIa, so disagreement with the two-process model is not too surprising, but we do not know why the sign of the discrepancy is different in these two cases. The intermediate mass stars that dominate AGB yields have lifetimes that are shorter than the 1 Gyr median delay of the SNIa enrichment, so the steeper trend of C+N is arguably the more natural result because of the larger change in C+N over the restricted time range of the thick disk compared to the interval between the thick and thin disks. Like Ce, Na shows a lower observed slope than predicted slope, though this difference is slight for Na. Within the high-Ia population, Na and Ce both show enhanced residual abundances at young ages W22, similar to the trends for the neutron capture elements Y and Ba in GALAH (Griffith et al., 2022), and they show similar structure of residual correlations with Galactic position (Griffith et al., 2024b). The physical origin of a link between Na and Ce enrichment is unclear, but our results provide further evidence that it exists. Ni also has a slightly steeper observed than predicted trend, which suggests that the Ni/Fe yield ratio of SNIa may be slightly higher in the thick disk population.
@@
The scatter about the [X/Mg]med vs. trend is higher for most elements than we predict based on the two-process model plus Gaussian scatter with the reported measurement error. In some cases, this excess scatter may indicate that the measurement uncertainties are underestimated, or at least that there are non-Gaussian tails to the error distribution. Recent studies have been mixed on how accurate the ASPCAP uncertainties are: Mead et al. (2024) found that the uncertainties are underestimated at dwarf galaxy metallicities while Sinha et al. (2024) found that the uncertainties were sufficiently well-estimated in open clusters near solar metallicity. The plateau sample analyzed in this work falls in between these two metallicity ranges. Future work on calibrating ASPCAP abundance uncertainties with high-resolution spectra will shed more light on this issue. Accurate abundance measurement uncertainties are essential for meaningful work on investigating the intrinsic abundance scatter of any population, not just the low-Ia plateau.
@@
Much of the attention in MW chemical evolution modeling has gone to understanding the difference in abundance trends between the thin and thick disks, and the bimodality of the [/Fe] distribution at a given [Fe/H]. These features remain incompletely understood, with a variety of possible explanations. Extending earlier studies of the [/Fe] scatter within the thick disk (Bertran de Lis et al., 2016; Vincenzo et al., 2021), our results show that there are significant variations of the SNIa/CCSN ratio within the low-Ia population at fixed [Mg/H], producing coherent variations in the abundance ratios of many elements. They further show distinctive behavior of some elements expected to have delayed contribution from AGB stars. Asteroseismic ages of APOGEE stars in the Kepler field show that the thick disk stars have a remarkably small age range, at least at the solar annulus (Pinsonneault et al., 2024). Producing the observed abundance variations within a short span of time is a challenge for models of the early evolution of the MW disk.
@@ 6 Acknowledgments
We thank Chris Hayes for providing upper limit relations and helpful discussion, Sten Hasselquist for detailed comments that helped improve the paper, and Yuan-Sen Ting for interesting discussion on the two-process model.

### v2

@@ 4.2 Scatter Around the SNIa/CCSN Correlation
Broadly, most of the more well-measured elements (e.g., Si, Ca, Al, Fe, Ni, Mn) show similar RSE between the data and prediction, indicating that measurement noise is sufficient to explain the remaining variation after contributions from a varying SNIa/CCSN ratio are removed. In other words, the amount of SNIa contribution and observational errors are the primary sources of variation on the plateau for these elements. Many of the elements that have larger observational uncertainties (e.g., S, Na, K, Cr, V, Co, Cu, and Ce) have higher observed RSE than predicted RSE. The explanation may be that the reported measurement uncertainties are still underestimated, or that the errors have non-Gaussian tails, since several of these elements (most notably S, Na, V, Cu, and Ce) have relatively weak lines that are often affected by blending. In these cases, the shallower spectral features make abundances difficult to measure and less reliable, particularly at lower abundances, which may increase the scatter. In some cases, especially Ce and perhaps C+N and Na, the larger observed RSE could reflect intrinsic scatter in an enrichment process that is neither CCSN nor SNIa.
@@ 4.3 Covariances Between Elements
As we are calculating the covariances of precomputed residuals, the expectation value in Equation 10, , is the mean of the residuals over the entire plateau sample which should be close to, but may not be exactly, 0. As the covariances are calculated for element pairs, the exclusion of any one element does not affect the presence of correlations (or lack thereof) between the remaining elements; similarly, the inclusion of a new element only adds the potential for correlations with that new element without affecting the others.
Figure 10 shows the covariance matrices for the observed residuals from the median, [X/Mg]med,obs, in the left panel and the covariance of residuals from the two-process model, [X/Mg]2proc, in the right panel. The size of each dot corresponds to the magnitude of the covariance, and the diagonals represent the variance of the residuals in each element. Blue filled dots represent positive covariances while open red dots represent negative covariances.
@@
The left panel of Figure 10 has many positive covariances, indicating that most elements tend to deviate from the median in the same direction. In particular, the Fe-peak elements, highlighted by the yellow square, have significant positive residuals; this is consistent with our previous result that variation in SNIa/CCSN contribution is a significant source of correlated abundance scatter on the low-Ia plateau. The broadly positive residuals within an element group are also seen, though to a lesser extent, in the elements highlighted by the purple box. We excluded the “Fe-cliff” element Cu from the Fe-peak yellow square here because it behaves differently from the other Fe-peak elements: Cu has (small) negative covariances with Cr and Co. Interestingly, Kobayashi et al. (2020a) find that Co and Cu are both largely produced by hypernovae, so one might expect these two elements in particular to be correlated.
@@
The two-process model captures the variation in the SNIa/CCSN ratio, so that the covariance in the two-process residuals should have the covariance from SNIa/CCSN variation removed. The right panel of Figure 10 shows the covariance matrix of the two-process residuals. As expected, the signs of the two-process residual covariances are much more mixed, compared to the mostly-positive median residual covariances. The magnitude of the covariances between Fe-peak elements have significantly decreased in the two-process residuals, and some have even changed sign, compared to the median residuals. However, V, Mn, and Co, which are odd-Z Fe-peak elements, still show some positive correlation with each other in the two-process residuals. The elements already have small covariances, and their magnitudes are not significantly different between the median and two-process residuals. However, the Ca-O and Ca-Si covariances have different signs. Similarly, the light odd-Z elements like C+N, Na, and Al do not show any obvious correlations with each other in their two-process residuals, though some covariance signs change from the median residuals. The two-process residuals of C+N are less correlated with those of the Fe-peak elements than the median residuals and are still positively correlated with Ce. Since C+N has contribution from a delayed process, this result indicates that this delayed process is not fully explained by the same delayed process (SNIa) producing the Fe-peak elements; furthermore, as Ce is an AGB element, this suggests that an AGB process also contributes to C+N.
@@
W22 removed outliers with a two-process residual greater than 10 times the measurement error for that element from their covariance calculations. If we also perform this cut, we find similar results to before, but the most affected element is Ce. More specifically, the Ce-(C+N) intrinsic [X/Mg]2proc covariance and Ce-Al [X/Mg]med,obs covariance change sign from positive to negative, while the Ce-Ni, Ce-Mn, and Ce-Co [X/Mg]med,obs covariances change from negative to positive. The rms scatter of Ce (diagonal element) also decreases. The majority of the removed outliers appear to be the same population of stars with strongly enhanced Ce and C+N found in W22, who propose that they are rare products of AGB mass transfer, where extra C+N and Ce from the AGB star are deposited onto the observed star’s photosphere. Their presence here indicates that such mass transfer events occur in stars on the low-Ia plateau.
@@ 5 Discussion
First, we tested the effect of the assumed Fe-to-Mg ratio from pure CCSN enrichment, [Fe/Mg]cc, on the two-process model as formalized by W22. When the process amplitudes and are inferred from only [Mg/H] and [Fe/Mg], the predicted linear {X/Mg} abundance is simply a linear interpolation (or extrapolation) of the median {X/Mg} ratios in the high- and low-Ia populations along {Fe/Mg}. It directly follows that the predicted abundances are independent of the assumed [Fe/Mg]cc, and we show that this independence holds to an excellent approximation even when using a multi-element fit for and .
@@
Our results clearly establish that there is a systematic difference in the SNIa/CCSN enrichment ratio between stars at the “top” of the low-Ia plateau and stars at the “bottom,” a range of roughly 0.1 dex in [Mg/Fe]. This rules out the idea that the median [Mg/Fe] on the low-Ia plateau represents the core collapse yield ratio [Mg/Fe]cc, with abundance ratios scattering stochastically above and below because of, e.g., stochastic sampling of the CCSN population. It remains possible that [Mg/Fe]cc lies at the upper boundary of the plateau, at [Mg/Fe] (See Figure 1). However, our results also lend credence to the idea that the true CCSN ratio is substantially higher, perhaps [Mg/Fe], and the observed low-Ia plateau already includes substantial SNIa enrichment (Maoz & Graur, 2017; Conroy et al., 2022). Maintaining a nearly flat plateau at this intermediate [Mg/Fe] requires accelerating star formation if the ISM is well mixed (Conroy et al., 2022; Chen et al., 2023), though a similar effect may be achieved if early enrichment is spread throughout the circumgalactic medium and only slowly returns to the star-forming ISM (Mason et al., 2024).
@@
For most elements we find good agreement between the predicted and observed slope of [X/Mg]med vs. , but there are exceptions. Most notably, the observed slope is steeper than predicted for C+N and shallower than predicted for Ce. For both abundances, we expect the delayed contribution comes from AGB stars rather than SNIa, so disagreement with the two-process model is not too surprising, but we do not know why the sign of the discrepancy is different in these two cases. The intermediate mass stars that dominate AGB yields have lifetimes that are typically shorter than the 1 Gyr median delay of the SNIa enrichment, which can vary from 0.18 to 2.5 Gyr depending on the model (see Table 2 of Dubay et al. 2024).333Dubay et al. (2024) find that delay time distributions with fewer prompt SNIa, which have have median delay times of 12.5 Gyr, result in multi-zone chemical evolution models that better fit the [/Fe] distribution in the MW disk. The steeper trend of C+N is arguably the more natural result because of the larger change in C+N over the restricted time range of the thick disk compared to the interval between the thick and thin disks. Ni also has a slightly steeper observed than predicted trend, which suggests that the Ni/Fe yield ratio of SNIa may be slightly higher in the low-Ia population.
@@
The scatter about the [X/Mg]med vs. trend is higher for most elements than we predict based on the two-process model plus Gaussian scatter with the reported measurement error. In some cases, this excess scatter may indicate that the measurement uncertainties are underestimated, or at least that there are non-Gaussian tails to the error distribution. Recent studies have been mixed on how accurate the ASPCAP uncertainties are for DR17: Mead et al. (2024) found that the uncertainties are underestimated at dwarf galaxy metallicities, while Sinha et al. (2024) found that the uncertainties were sufficiently well-estimated in open clusters near solar metallicity. The plateau sample analyzed in this work falls in between these two metallicity ranges. Future work on calibrating ASPCAP abundance uncertainties with high-resolution spectra will shed more light on this issue. Accurate abundance measurement uncertainties are essential for meaningful work on investigating the intrinsic abundance scatter of any population, not just the low-Ia plateau. Nonetheless, our main result that differences in the SNIa/CCSN ratio are a significant cause of variation on the low-Ia plateau holds regardless of potential changes to the magnitude or distribution of measurement errors because it is based on the slope of the [X/Mg]med vs. trend, not the scatter around it.
@@
Much of the attention in MW chemical evolution modeling has gone to understanding the difference in abundance trends between the thin and thick disks, and the bimodality of the [/Fe] distribution at a given [Fe/H]. These features remain incompletely understood, with a variety of possible explanations. Extending earlier studies of the [/Fe] scatter within the low-Ia population (Bertran de Lis et al., 2016; Vincenzo et al., 2021), our results show that there are significant variations of the SNIa/CCSN ratio within the low-Ia population at fixed [Mg/H], producing coherent variations in the abundance ratios of many elements. However, the source of the variation itself is not obvious; potential causes may include stellar migration and/or uneven mixing of the ISM during the formation of the plateau. They further show distinctive behavior of some elements expected to have delayed contribution from AGB stars. Asteroseismic ages of APOGEE stars in the Kepler field show that the thick disk stars have a remarkably small age range, at least at the solar annulus (Pinsonneault et al., 2024). Producing the observed abundance variations within a short span of time is a challenge for models of the early evolution of the MW disk.
@@ 6 Acknowledgments
We thank Chris Hayes for providing upper limit relations and helpful discussion, Sten Hasselquist for detailed comments that helped improve the paper, and Yuan-Sen Ting for interesting discussion on the two-process model. We also thank the anonymous referee for helpful feedback.
@@ Appendix A Upper Limits
Appendix B Reproducibility
We have produced a public Git repository containing a Jupyter notebook that reproduces all figures in this paper. Included with the repository are also data tables for values assuming different [Fe/Mg]pl (for reproducing Figure 2) and upper limits of the plateau sample in all elements (for reproducing Figure 11). The repository is accessible at https://github.com/tawnysit/plateau_variation.